OpenStudy (dancing_stars23):

medal if you check A pizza place offers 5 different cheeses and 10 different toppings. In how many ways can a pizza be made with 2 cheese and 3 toppings? 1200 1000 800 600 so I know you go 5!/2!*3! = 10 10!/3!*7! =120 would the answer be 1200 right? I have never done this before we two things before. just want to make sure!

OpenStudy (mathstudent55):

Since order does not matter, this is a combination problem. To pick 2 cheeses out of 5, you have \(_5C_2\) To pick 3 toppings out of 10, you need \(_{10}C_3\) To find the total number, multiply those results. \(_5C_2 \times _{10}C_3 = \dfrac{5!}{(5 - 2)!2!} \times \dfrac{10!}{(10 - 3)!3!} \) \(= \dfrac{5!}{3!2!} \times \dfrac{10!}{7!3!} = \dfrac{5 \times 4}{2} \times \dfrac{10 \times 9 \times 8}{3 \times 2}\) \(= 10 \times 120 = 1200\)

OpenStudy (mathstudent55):

You are correct.

OpenStudy (dancing_stars23):

@mathstudent55 thanks so much!

OpenStudy (mathstudent55):

You're welcome.